OpenStudy (anonymous):

HELP!!!!!! Prove by mathematical induction that a^n - b^n is divisible by (a-b) for n, all natural numbers

OpenStudy (anonymous):

Induction always has 3 steps. First check for n=1 or sometimes n=2. (first and/or second step/s) So here for n=1 we have a-b which is divisible by a-b for n=2 we have a^2-b^2=(a-b)(a+b) so divisible by a-b Check!

OpenStudy (anonymous):

Now for the second step we assume that our claim is correct and we will use this to try and solve for n+1. So assumtion: a^n - b^n is divisible by (a-b) for n

OpenStudy (anonymous):

Last step where we need to calculate is to prove for n+1

OpenStudy (anonymous):

Yeah that's my problem the calculation step for n+1

OpenStudy (kirbykirby):

I think I found a solution to this, that requires a not-so obvious step. You need to show that \(a^{n+1}-b^{n+1}\) is dividible by \(a-b\), or in other words: You need to show \[\frac{a^{n+1}-b^{n+1}}{a-b} \] results in an integer. So: \[\begin{align}\frac{a^{n+1}-b^{n+1}}{a-b}&=\frac{a^na-b^nb}{a-b}\\~ \\&=\frac{a^na-b^nb\color{red}{+ab^n-ab^n}}{a-b}\\~ \\&=\frac{a(a^n-b^n)+b^n(-b+a)}{a-b} ,\text{factor by combining } \\&~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\text{1st-4th terms, 2nd and 3rd terms}\\&=\frac{a(a^n-b^n)}{a-b}+b^n\end{align} \] By the induction hypothesis, \(\dfrac{a^n-n^n}{a-b}\) results in an integer, say \(p\), so \(ap+b^n\) is an integer, and so the induction is done.

OpenStudy (anonymous):

THANK YOU

OpenStudy (kirbykirby):

=] yw